OpenStudy (bluestar70556):

Evaluate: The Limit of x to infinity 2^x + x^3/x^2 + 3^x My choices are 0, 1, 3/2, 2/3, and The limit does not exist. However I'm pretty sure that the answer is 2/3

OpenStudy (dallascowboys88):

and btw i think you're right about your answer :)

OpenStudy (bluestar70556):

Okay, Thank you! :) I just wanted to make sure

OpenStudy (eliesaab):

No that is not the right answer

OpenStudy (bluestar70556):

Okay, How would I calculate the correct answer?

OpenStudy (eliesaab):

Let me convince you that the limit is 0

OpenStudy (bluestar70556):

The only way I could see the limit being 0 is if we look at the bases

OpenStudy (eliesaab):

\[ 2^x + x^3\approx 2^x \] when x is near \( \infty\)

OpenStudy (eliesaab):

\[ x^2 + 3^x\approx 3^x \] when x is near \( \infty\)

OpenStudy (eliesaab):

The ratio is like \[ \frac{2^x}{3^x}= \left ( \frac 2 3 \right)^x \to 0\] when x goes to \(\infty \)

OpenStudy (eliesaab):

Exponential growth is the deciding part near \( \infty \)

OpenStudy (bluestar70556):

Ohhhh Okay, and in my lecture they looked at bases when it came to exponential functions. So because 3 is the larger base and is in the denominator, it automatically goes to 0, I see now. Thank you for explaining that!!

OpenStudy (eliesaab):

A more formal proof goes as \[ \frac{x^3+2^x}{x^2+3^x}=\frac{x^3+2^x}{\frac{3^x \left(x^2+3^x\right)}{3^x}}=\frac{\frac{x^3}{3^x}+\left(\frac{2}{3}\right)^ x}{\frac{x^2}{3^x}+1}\] and let x goes to \(\infty\)

OpenStudy (eliesaab):

I divided up and down by \[ 3^x \]

OpenStudy (eliesaab):

YW